# On the Consistency of Multilingual Context Utilization in Retrieval-Augmented Generation

### v2

@@ 4.2 Interpretability-based Assessment
Dataset In-lang Out-lang
CTI Thres = AVG + 1.0 * SD
XQUAD 98 99
MKQA 100 100
CTI Thres = AVG + 1.5 * SD
XQUAD 97 99
MKQA 100 100
CTI Thres = AVG + 2.0 * SD
XQUAD 95 98
MKQA 100 100
Table 2: Percentage (%) of context-sensitive responses when Aya is provided with in-language vs. out-language gold passages, detected by MIRAGE under different CTI thresholds.
To further verify our findings that the passage language is not a barrier to LLMs’ understanding capability of the multilingual retrieved passages, we adopt MIRAGE (Qi et al., 2024), a model internal-based method for attributing model responses to the retrieved passages in RAG systems. Generally, it consists of two components: (1) CTI for detecting contextual sensitivity for the generated sentence and (2) CCI for attributing the detected sentences back to each retrieved passage. Given the single-passage setup, in this section we only use the CTI module for evaluating the passage dependency of the model response. For each generated token, this module measures the shift in output probability distribution when no context vs. one passage is provided, measured by KL divergence (Kullback & Leibler, 1951), while keeping the generated sentence prefix fixed. If at least one token is higher than an empirically set CTI threshold, the generated sentence is marked as sensitive to the context provided in the prompt.
[…]
5 Multi-Passage mRAG
5.1 Accuracy Results
@@
Table 3 presents the results, including the no-context baseline and single in-/out-language gold passage results as computed in Section 4, to enable comparison (see Appendix F for full language-specific results). For this analysis, we also consider as valid the questions that were answered correctly but in the wrong language, as they also reflect a proper understanding of the context by the model. Interestingly, models provided with 3 out-language gold passages achieve higher accuracy than when provided with a single in-language gold passage in the query language, emphasizing the potential of cross-lingual retrieval for mRAG. As expected, the presence of distractors leads to lower accuracy. Notably, this is true for all models, datasets, and setups. However, the effect is considerably stronger in MKQA than in GMMLU-Choice, likely due to the stricter lexical-matching metric adopted for MKQA. We also verify that a higher proportion of distractors (3/4 vs. 1/4) is much more harmful for answer accuracy, which confirms the importance of having access to a high-quality cross-lingual retriever (Chirkova et al., 2024). When comparing the drop between in-language distractors and out-language distractors, we find that in-language distractors have a larger impact in most cases, matching our hypothesis that this is a particularly challenging scenario for LLMs. However, differences are small in many cases, indicating the language of the distractor is not a major issue for multi-passage mRAG.
@@ 3 Gold (out) + 1 Distractor
# Contextual 0.85 0.79 0.92 0.89
# Influential 0.35 0.25 0.50 0.43
Table 4: Average number of distractors containing contextual cues (# Contextual) and receiving a higher sum of CCI scores than all gold passages (# Influential), for Aya.
@@ 5.2 Interpretability-based Evaluation
We adopt once again MIRAGE (Qi et al., 2024) to understand how the internal model dynamics are affected by our various simulated multi-passage mRAG scenarios. We sample 50 instances from each dataset and use MIRAGE to attribute Aya-expanse-8B responses to the provided passages via contrastive feature attribution (Yin & Neubig, 2022). Then, we compute # Contextual: the average number of distracting passages that contain at least one contextual cue for the produced answer (i.e. a token marked by CCI in MIRAGE), and # Influential: the average number of distractors that receive a higher sum of CCI attribution scores than all gold passages for each query.
@@ 6 Conclusion
At the same time, we also detected some important directions for future improvement. Firstly, poor generation abilities in many languages push the models to respond in a different language than that of the query, resulting in answers that would be deemed useless by most end-users. Importantly, we showed that this also happens when the retrieval works optimally. This suggests that, rather than just trying to optimize the retriever, it may be more effective to invest on the model generation abilities in a specific (set of) user language(s) –for instance by continued pre-training (Fujii et al., 2024; Gao et al., 2024) on generic corpora of those languages– or to apply techniques that push the model to decode in a given language, such as contrastive decoding (Li et al., 2023; O’Brien & Lewis, 2023). Secondly, the presence of distracting passages (i.e., relevant to the query topic, but not directly functional to answer it) in the context can have a very negative effect on answer accuracy in open-domain QA. While this effect is rather similar regardless of the distractors’ language, it does highlight the importance of carefully ranking the retrieved passages and to aim for precision when selecting which passages are provided to the model.
The limitations of our work include relying on a strict lexical matching of the answer to compute model accuracy and to detect gold passages. While commonly used, this approach is sensitive to minor variations or rephrasings of the answers and led to a serious underestimation of model performance with out-language gold passages in one of our QA datasets, XQUAD. Additionally, this approach may overlook passages that provide valuable information but in a slightly rephrased form compared to the gold answer. On the retrieval side, simulating cross-language retrieval by combining results of N in-language retrievers may yield a more comprehensive set of passages than what we could obtain from a single run of a cross-language retriever. While this does not affect our results on the side of context utilization, it may overestimate retriever performance when our findings are applied to real-world mRAG systems. In terms of datasets, XQUAD was the only one including parallel gold passages, which allowed us to fully isolate the effect of a passage language from that of its content. However, its extractive QA nature makes it less representative of realistic mRAG tasks, highlighting the need to develop better parallel mRAG datasets in future work.
@@
To conclude, our work underscores the potential of cross-lingual retrieval in enhancing multilingual QA performance, and stresses the importance of focusing not only on retrieval optimization, but also on improving language-specific generation capabilities. We believe this dual focus will be key to unlocking more robust, user-friendly mRAG systems that can operate effectively across diverse language settings.
@@ Appendix A Dataset Examples
Appendix B Full Statistics of the Filtered MKQA and GMMLU Datasets
@@ Appendix C Subjects Covered by the Filtered GMMLU Set
Appendix D Prompts and Instructions
@@
Appendix E Error Analysis on XQUAD
{CJK*}
@@ Appendix E Error Analysis on XQUAD
UTF8gbsn While our MKQA and GMMLU results strongly suggest our studied LLMs can understand the provided passages regardless of their language, the in-/out-language gap in XQUAD remains unexplained. To address this, we conduct a manual error analysis on XQUAD with Aya-Expanse-8B, focusing on a random sample of 20 Spanish and 20 Chinese queries that were answered correctly when provided with in-language passages, but wrongly with out-language passages. In most cases, we observe that models successfully understood the context and generated a proper response, however, this response did not perfectly match the gold answer provided in the dataset. This can be due to the presence of synonyms or slight paraphrases of the gold answer, or –in the case of languages with different scripts– to transliteration variations (Knight & Graehl, 1997). For instance, the gold answer for a Spanish question is ‘evolución de la lengua y la literatura alemanas’ (i.e. ‘evolution of the German language and literature’). In the in-language setup, the model manages to generate this exact string as it is included in the provided Spanish passage. However, when the same passage is provided in English, the model generates the semantically equivalent phrase ‘… evolución del idioma y la literatura alemana…’, or ‘…desarrollo del idioma y la literatura alemana…’ when the passage is provided in Chinese. Similarly, for a Chinese query with gold answer ‘亚里士多德宇宙学’ (i.e. ‘Aristotelian cosmology’), model responses slightly differ when provided with different out-language passages (e.g. ‘亚里士多德宇宙论’, ‘阿里斯托的宇宙论’, or ‘阿里斯托特利宇宙论’ with English, Arabic, or Greek passage respectively), all of which are correct translations of ‘Aristotelian cosmology’. While this issue can always affect lexical-matching evaluation, it is particularly severe in XQUAD as many answers in this dataset are named entities or sentence segments due to the extractive nature of the task, which in turn causes an underestimation of the models capability.
Appendix F Language-specific Results
@@
The detailed results for each query-passage language pair on XQUAD are given in Figure 5. The detailed single-passage mRAG results for each language on all datasets are provided in Table 10 to 15. The detailed results for each language in the multi-passage mRAG experiments are shown in Table 16 to 19.

### v3

@@ 4.2 Interpretability-based Assessment
To further verify our findings that the passage language is not a barrier to LLMs’ understanding capability of the multilingual retrieved passages, we adopt MIRAGE (Qi et al., 2024), a model internal-based method for attributing model responses to the retrieved passages in RAG systems. Generally, it consists of two components: (1) CTI for detecting contextual sensitivity for the generated sentence and (2) CCI for attributing the detected sentences back to each retrieved passage. Given the single-passage setup, in this section we only use the CTI module for evaluating the passage dependency of the model response. For each generated token, this module measures the shift in output probability distribution when no context vs. one passage is provided, measured by KL divergence (Kullback and Leibler, 1951), while keeping the generated sentence prefix fixed. If at least one token is higher than an empirically set CTI threshold, the generated sentence is marked as sensitive to the context provided in the prompt.
Dataset AVG+1.0*SD AVG+1.5*SD AVG+2.0*SD
In. Out. In. Out. In. Out.
XQUAD 98 99 97 99 95 98
MKQA 100 100 100 100 100 100
Table 2: Percentage (%) of context-sensitive responses when Aya is provided with in-language (In.) vs. out-language (Out.) gold passages, detected by MIRAGE under different CTI thresholds.
@@
We select Aya-Expanse-8B as the studied model and sample 500 instances separately from XQUAD and MKQA. Table 2 shows the results under different CTI thresholds. We find that nearly all generated responses are tagged as context-sensitive by MIRAGE, even when setting a higher CTI threshold ($\rm avg$} + 2 $\rm std\_dev$}) than the one used in the original paper. This confirms that the provided passage significantly drives models’ predictions regardless of its language.
5 Multi-Passage mRAG
5.1 Accuracy Results
Table 3 presents the results, including the no-context baseline and single in-/out-language gold passage results as computed in Section 4, to enable comparison (see Appendix I for full language-specific results). For this analysis, we also consider as valid the questions that were answered correctly but in the wrong language, as they also reflect a proper understanding of the context by the model. Interestingly, models provided with 3 out-language gold passages achieve higher accuracy than when provided with a single in-language gold passage in the query language, emphasizing the potential of cross-lingual retrieval for mRAG. As expected, the presence of distractors leads to lower accuracy. Notably, this is true for all models, datasets, and setups. However, the effect is considerably stronger in MKQA than in GMMLU-Choice, likely due to the stricter lexical-matching metric adopted for MKQA. We also verify that a higher proportion of distractors (3/4 vs. 1/4) is much more harmful for answer accuracy, which confirms the importance of having access to a high-quality cross-lingual retriever (Chirkova et al., 2024). When comparing the drop between in-language distractors and out-language distractors, we find that in-language distractors have a larger impact in most cases, matching our hypothesis that this is a particularly challenging scenario for LLMs. However, differences are small in many cases, indicating the language of the distractor is not a major issue for multi-passage mRAG.
@@ 3 Gold (out) + 1 Distractor
# Context. 0.85 0.79 0.92 0.89
# Influent. 0.35 0.25 0.50 0.43
Table 4: Average number of distractors containing contextual cues (# Context.) and receiving a higher sum of CCI scores than all gold passages (# Influent.), for Aya.
@@ 5.2 Interpretability-based Evaluation
We adopt once again MIRAGE (Qi et al., 2024) to understand how the internal model dynamics are affected by our various simulated multi-passage mRAG scenarios. We sample 50 instances from each dataset and use MIRAGE to attribute Aya-expanse-8B responses to the provided passages via contrastive feature attribution (Yin and Neubig, 2022). Then, we compute # Contextual: the average number of distracting passages that contain at least one contextual cue for the produced answer (i.e. a token marked by CCI in MIRAGE), and # Influential: the average number of distractors that receive a higher sum of CCI attribution scores than all gold passages for each query.
@@ 6 Conclusion
At the same time, we also detected some important directions for future improvement. Firstly, poor generation abilities in many languages push the models to respond in a different language than that of the query, resulting in answers that would be deemed useless by most end-users. Importantly, we showed that this also happens when the retrieval works optimally. This suggests that, rather than just trying to optimize the retriever, it may be more effective to invest on the model generation abilities in a specific (set of) user language(s) –for instance by continued pre-training (Fujii et al., 2024; Gao et al., 2024) on generic corpora of those languages– or to apply techniques that push the model to decode in a given language, such as contrastive decoding (Li et al., 2023; O’Brien and Lewis, 2023). Secondly, the presence of distracting passages (i.e., relevant to the query topic, but not directly functional to answer it) in the context can have a very negative effect on answer accuracy in open-domain QA. While this effect is rather similar regardless of the distractors’ language, it does highlight the importance of carefully ranking the retrieved passages and to aim for precision when selecting which passages are provided to the model.
@@
To conclude, our work underscores the potential of cross-lingual retrieval in enhancing multilingual QA performance, and stresses the importance of focusing not only on retrieval optimization but also on improving language-specific generation. We believe this dual focus will be key to unlocking more robust and user-friendly mRAG systems that can operate effectively across diverse language settings.
Limitations
Our work relies on strict lexical matching to compute model answer accuracy and to detect gold passages. While commonly used, this approach is sensitive to minor variations or rephrasings of the answers and led to a serious underestimation of model performance with out-language gold passages in one of our QA datasets, XQUAD. In our paper, we have tested and reported BERTScore and LLM-based evaluation on XQUAD, as detailed in Section 3.3 to enhance the robustness of our findings. These semantic evaluations mirror the trends observed with the lexical metric, mitigating—if not eliminating—the risk that paraphrasing may influence the results. Nevertheless, future work could incorporate broader metrics and benchmarks to make the assessment more comprehensive.
Additionally, the use of lexical matching in detecting gold passages may overlook passages that provide valuable information but in a slightly rephrased form compared to the gold answer. Nonetheless, Table 3 shows that attaching even a single distracting passage identified by this heuristic method substantially degrades model accuracy. Thus, despite its limitations, lexical matching proves to be a practical and effective way for locating distracting passages in our experimental setting. Future work could explore more semantic retrieval methods to capture paraphrased gold evidence.
On the retrieval side, simulating cross-language retrieval by combining results of N in-language retrievers may yield a more comprehensive set of passages than what we could obtain from a single run of a cross-language retriever. While this does not affect our results on the side of context utilization, it may overestimate retriever performance when our findings are applied to real-world mRAG systems. In terms of datasets, XQUAD was the only one including parallel gold passages, which allowed us to fully isolate the effect of a passage language from that of its content. However, its extractive QA nature makes it less representative of realistic mRAG tasks, highlighting the need to develop better parallel mRAG datasets in future work.
@@ Appendix A Dataset Examples
Appendix B Full Statistics of the Filtered MKQA and GMMLU Datasets
@@ Appendix C Subjects Covered by the Filtered GMMLU Set
Appendix D Extensive Evaluation
Lang. BERTScore LLM-Based Score
IN. OUT. IN. OUT.
en 90.27 82.58 93.19 81.54 (+12.93)
ar 82.87 81.55 91.93 63.98 (+23.25)
de 81.60 80.69 90.25 71.73 (+14.86)
el 82.38 81.24 92.18 66.00 (+21.04)
es 82.15 81.14 94.37 73.05 (+14.62)
hi 83.31 82.15 89.75 65.78 (+16.71)
ro 81.68 80.65 91.93 69.11 (+18.69)
ru 83.22 81.95 91.51 66.78 (+21.18)
th 84.21 83.23 86.81 57.78 (+23.02)
tr 81.31 80.18 88.40 63.39 (+23.37)
vi 82.81 81.62 89.75 65.81 (+23.05)
zh 84.07 82.95 90.34 65.39 (+21.46)
Table 8: BERTScore (F1) and LLM-based evaluation (Accuracy) on XQUAD with AYA. The numbers between brackets indicate the proportion of queries that are correctly answered but in the wrong language.
Since exact matching could be overly strict for the evaluation, we further adopt two complementary metrics on XQUAD with AYA.
Semantic similarity (BERTScore)
We compute BERTScore, serving as a language-agnostic metric, between each model response and its ground-truth answer based on the semantic similarity of model responses with the gold answer. Table 8 shows that models achieve comparable F1 scores in all query languages when fed gold passages in- or out-language. This finding is in line with our claim that LLMs are capable of understanding the gold passages regardless of their languages.
LLM-based evaluation (GPT-4.1-nano)
However, semantic similarity cannot capture language mismatching. Therefore, we prompted GPT-4.1-nano to judge whether each response matches (i) the correct answer and (ii) its translation in the passage language. As shown in Table 8, overall accuracy on board is higher than lexical-matching accuracy in our paper, but the trend remains: models score better on IN than on OUT. If we allow “correct answer in the wrong language” as acceptable, the IN/OUT gap almost disappears.
Taken together, both semantic and LLM-based evaluation support our claim that LLMs are able to understand the multilingual gold passages regardless of their languages, but suffer from decoding the answer correctly in the user query language.
Appendix E Prompts and Instructions
@@
Appendix F Error Analysis on XQUAD
@@
While our MKQA and GMMLU results strongly suggest our studied LLMs can understand the provided passages regardless of their language, the in-/out-language gap in XQUAD remains unexplained. To address this, we conduct a manual error analysis on XQUAD with Aya-Expanse-8B, focusing on a random sample of 20 Spanish and 20 Chinese queries that were answered correctly when provided with in-language passages, but wrongly with out-language passages. In most cases, we observe that models successfully understood the context and generated a proper response, however, this response did not perfectly match the gold answer provided in the dataset. This can be due to the presence of synonyms or slight paraphrases of the gold answer, or –in the case of languages with different scripts– to transliteration variations (Knight and Graehl, 1997). For instance, the gold answer for a Spanish question is ‘evolución de la lengua y la literatura alemanas’ (i.e. ‘evolution of the German language and literature’). In the in-language setup, the model manages to generate this exact string as it is included in the provided Spanish passage. However, when the same passage is provided in English, the model generates the semantically equivalent phrase ‘… evolución del idioma y la literatura alemana…’, or ‘…desarrollo del idioma y la literatura alemana…’ when the passage is provided in Chinese. Similarly, for a Chinese query with gold answer ‘亚里士多德宇宙学’ (i.e. ‘Aristotelian cosmology’), model responses slightly differ when provided with different out-language passages (e.g. ‘亚里士多德宇宙论’, ‘阿里斯托的宇宙论’, or ‘阿里斯托特利宇宙论’ with English, Arabic, or Greek passage respectively), all of which are correct translations of ‘Aristotelian cosmology’. While this issue can always affect lexical-matching evaluation, it is particularly severe in XQUAD as many answers in this dataset are named entities or sentence segments due to the extractive nature of the task, which in turn causes an underestimation of the models capability.
Appendix G Advanced System Prompting
In our main experiments, we follow the previous works (Chirkova et al., 2024; Zhang et al., 2024) and adopt the direct prompt. To test if a stronger prompt could mitigate language-mismatch errors, we add a two-step instruction that first allows the model to answer in any appropriate language, then explicitly translates the answer into the query language. Formally: ‘Write a high-quality answer to the given question using the provided search results. Please respond in English. Specifically, please follow the two steps below. Step 1: Generate a complete answer to the question in any appropriate language. Step 2: Translate your entire answer into clear, natural‐sounding English.’
Same as the main experiment in the paper, the prompt is translated into other query languages and explicitly specifies the desired generation language. For instance, the prompt for Spanish queries is:
‘Escriba una respuesta de alta calidad a la pregunta dada utilizando los resultados de búsqueda proporcionados. Por favor responda en español. Específicamente, siga los dos pasos a continuación. Paso 1: Genere una respuesta completa a la pregunta en cualquier idioma apropiado. Paso 2: Traduce toda tu respuesta a un español claro y con sonido natural.’
Language LLM-Based Score
IN. OUT.
en 94.71 82.52 (+11.24)
ar 87.14 64.38 (+24.85)
de 89.92 73.12 (+13.76)
el 90.59 65.11 (+22.05)
es 92.61 73.51 (+14.23)
hi 85.13 63.80 (+20.06)
ro 92.61 67.00 (+19.17)
ru 90.67 69.49 (+19.34)
th 86.89 60.04 (+25.12)
tr 85.46 63.46 (+21.35)
vi 91.34 67.16 (+22.85)
zh 90.84 69.60 (+20.28)
Table 11: LLM-based evaluation score (Accuracy) on XQUAD with Aya, where stronger multi-step reasoning prompts are adopted. Nonetheless, the language-mismatching issue persists.
We run this prompt on XQUAD with AYA and evaluate via GPT-4.1-nano, the same setups and LLM-based evaluation as above in Appendix D. As shown in Table 11, compared to the original prompts, these stronger instructions reduced, but did not eliminate, the gap between in-language and out-language accuracy. Specifically, many responses still contained correct answers but remained in the wrong passage language, indicating that even explicitly guiding the LLM to do ‘think then translation’ cannot fully resolve decoding failures. These results underscore that decoding, rather than understanding, remains a substantial bottleneck.
Language Accuracy (Gemma3-27B-IT)
Non. In. Out. (AVG) Out. (Best) Out. (Worst)
en 21.3 86.7 67.1 (+4.9) 72.9 (+5.5) 60.3 (+9.4)
es 16.6 72.9 53.2 (+5.8) 59.2 (+9.4) 48.2 (+7.6)
de 17.0 72.5 52.0 (+4.7) 56.9 (+7.2) 48.2 (+7.0)
ro 14.5 76.2 52.0 (+4.3) 56.6 (+8.3) 47.4 (+6.2)
vi 15.1 77.6 49.9 (+7.4) 53.7 (+13.4) 44.3 (+9.7)
tr 12.2 67.4 45.4 (+5.9) 50.5 (+9.8) 42.1 (+5.0)
el 10.4 68.9 40.9 (+8.3) 44.7 (+16.1) 38.2 (+6.8)
zh 13.9 79.2 44.1 (+13.7) 45.7 (+20.7) 41.5 (+7.3)
ar 8.8 65.8 35.5 (+9.5) 37.4 (+19.6) 32.3 (+5.8)
hi 13.3 74.8 42.6 (+6.0) 46.8 (+15.0) 38.7 (+4.2)
ru 11.4 66.6 40.6 (+9.7) 42.6 (+11.8) 36.9 (+6.7)
th 11.1 74.6 35.4 (+14.2) 36.4 (+30.0) 34.5 (+9.8)
Language Accuracy (GPT5-nano)
Non. In. Out. (AVG) Out. (Best) Out. (Worst)
en 25.6 74.0 55.1 (+7.1) 58.3 (+8.3) 51.8 (+7.8)
es 20.2 64.8 50.2 (+3.6) 55.1 (+7.2) 46.3 (+2.4)
de 20.1 60.2 48.0 (+1.9) 52.0 (+6.2) 44.9 (+0.8)
ro 19.2 59.8 46.1 (+2.6) 49.4 (+7.3) 42.9 (+1.4)
vi 18.3 58.8 45.1 (+4.4) 47.8 (+5.0) 42.4 (+2.1)
tr 15.5 49.2 38.0 (+5.6) 39.8 (+7.4) 35.3 (+2.4)
el 12.5 55.5 35.6 (+6.1) 39.3 (+16.2) 32.9 (+2.1)
zh 16.5 59.6 39.1 (+8.9) 41.4 (+18.2) 37.8 (+3.6)
ar 10.3 48.2 31.3 (+8.8) 33.1 (+7.1) 30.0 (+3.9)
hi 12.8 45.2 29.8 (+10.3) 32.3 (+22.6) 27.0 (+13.3)
ru 12.8 48.9 34.1 (+6.4) 36.4 (+14.0) 31.8 (+3.9)
th 11.9 53.4 32.6 (+10.6) 34.2 (+25.2) 30.4 (+2.1)
Table 12: Language-specific results on XQUAD with larger LLMs (Gemma3-27B-IT and GPT5-nano). Numbers between brackets indicate the proportion of queries that are correctly answered but in the wrong language (i.e., not the query language).
Appendix H Extended Evaluation on Larger Models
To enhance the robustness of our experiments, we repeat the XQuAD evaluation (using the same setup) on a 27B open-source model (Gemma3-27B-IT) and a closed-source model estimated at 8-18B parameters (GPT5-nano) [^9] https://www.r-bloggers.com/2025/08/how-many-parameters-does-gpt-5-have/. The results in Table 12 show that, although overall accuracy of out-language passages improves, it remains substantially lower than on in-language passages. Moreover, a non-negligible fraction of questions are answered correctly in content but produced in the wrong language when the model receives out-language passages. These findings are consistent with Section 4.1 and further strengthen the generalization of our findings.
Appendix I Language-specific Results
@@
The detailed results for each query-passage language pair on XQUAD are given in Figure 4. The detailed single-passage mRAG results for each language on all datasets are provided in Table 13 to 18. The detailed results for each language in the multi-passage mRAG experiments are shown in Table 19 to 22.